# Hydrodynamic flows induced by localized torques (rotlets) in wedge-shaped geometries

### v1

@@ IV.2 Transverse torque
These results yield expressions for the velocity field in full agreement with those reported by Spagnolie and Lauga [17].
@@ V Hydrodynamic mobilities
The solution for the flow field derived above can be used to evaluate how the confining wedge influences the motion of a small spherical colloidal particle of radius $a$ subjected to an external torque. Due to hydrodynamic interactions with the no-slip boundaries, the particle will in general be displaced, and its rotational velocity will be affected. The leading-order term in the hydrodynamic coupling mobility is obtained by evaluating the image solution for the induced flow field at the position of the particle. This procedure yields an expression for the mobilities in terms of the small parameter $\epsilon=a/d$, where $d=\rho\sin(\alpha-\beta)$ and $a$ is the particle radius. The change in rotational mobility when compared to the situation in an unbounded fluid is obtained by evaluating half the vorticity of the flow resulting from the image solution at the position of the rotating particle.
@@
Defining $\mu_{0}^{rr}=1/(8\pi\eta a^{3})$ as the bulk hydrodynamic rotational mobility in an unbounded fluid, the translational and rotational velocities resulting from the hydrodynamic interactions with the boundaries are related to the hydrodynamic torque $\bm{L}$ acting on the particle through
@@
$\bm{V}=\left.\bm{v}(\bm{r})-\bm{v}^{\infty}(\bm{r})\right|_{\bm{r}=\bm{r}_{0}}\,,\qquad\bm{\Omega}=\left.\mu_{0}^{rr}\bm{L}+\tfrac{1}{2}\,\bm{\nabla}\times\left(\bm{v}(\bm{r})-\bm{v}^{\infty}(\bm{r})\right)\right|_{\bm{r}=\bm{r}_{0}}\,.$ (65)
Accordingly,
$\bm{V}=-\mu_{0}^{tr}\,\epsilon^{2}\bm{A}\cdot\bm{L}\,,\qquad\bm{\Omega}=\mu_{0}^{rr}\left(\bm{I}-\epsilon^{3}\bm{B}\right)\cdot\bm{L}\,$, (66)
where $\mu_{0}^{tr}=1/(6\pi\eta a^{2})$ is used to nondimensionalize the translation–rotation coupling, $\bm{I}$ denotes the identity matrix, and
$\bm{A}=\begin{pmatrix}0&0&a_{rz}\\ 0&0&a_{\theta z}\\ a_{zr}&a_{z\theta}&0\end{pmatrix},\qquad\bm{B}=\begin{pmatrix}b_{rr}&b_{r\theta}&0\\ b_{\theta r}&b_{\theta\theta}&0\\ 0&0&b_{zz}\end{pmatrix}$ (67)
@@
are dimensionless matrices, corresponding to the rotational–translational coupling mobilities and purely rotational mobilities, respectively. It follows that, to leading order, the coupling mobilities scale as $\epsilon^{2}$, while the correction to the rotational mobilities exhibit an even weaker scaling of $\epsilon^{3}$. By the symmetry of the mobility tensor, it follows that $b_{r\theta}=b_{\theta r}$. Accordingly, we are left with determining eight components that fully characterize the response to the applied torque.
We mention that for $\bm{A}$, related to rotational–translational coupling in the hydrodynamic mobilities, the prefactor is $6\pi\eta a^{2}$. Conversely, for the resistance it should be $8\pi\eta a^{2}$, as used in Sano and Hasimoto [9]. As a side remark, to compare with the dimensionless matrix $\bm{t}$ associated with rotational–translational coupling as defined in Ref. 9, we can note the relation $\bm{A}=\bm{t}^{\top}$. Generally, the grand mobility tensor and the grand resistance tensor are inverses of each other. The following results are derived using Eq. (65) together with the solution for the flow field as given by Eq. (52).
@@ V.1 Hydrodynamic coupling mobilities
Figure 5: Variations of the leading-order corrections to the different components of the hydrodynamic coupling mobilities as functions of $\beta/\alpha$ for various values of the semi-opening angle $\alpha$ of the wedge. To leading order, the corrections vanish for $\alpha=\pi/2$ in panels (a) and (c), and for both $\alpha=0$ and $\alpha=\pi/2$ in panels (b) and (d). Some curves are omitted to avoid overcrowding the figure.
@@
$a_{iz}=\frac{3}{64}\,\sin^{2}(\alpha-\beta)\int_{0}^{\infty}\left(M_{i}\operatorname{ch}(\beta\nu)+N_{i}\operatorname{sh}(\beta\nu)\right)\operatorname{sch}(\pi\nu)\,\mathrm{d}\nu\,$ (68)
@@
$M_{r}$ $=\left(12\Lambda_{x}^{\dagger}-\left(4\nu^{2}+7\right)H_{x}^{\dagger}\right)\cos\beta+\left(12\Lambda_{y}^{\dagger}-\left(4\nu^{2}+7\right)H_{y}^{\dagger}\right)\sin\beta+4\Lambda_{w}^{\dagger}\,$, (69a)
$N_{r}$ $=\left(12\Lambda_{x}-\left(4\nu^{2}+7\right)H_{x}\right)\cos\beta+\left(12\Lambda_{y}-\left(4\nu^{2}+7\right)H_{y}\right)\sin\beta+4\Lambda_{w}\,$, (69b)
@@
$M_{\theta}$ $=4\left(\left(\left(H_{x}-2\Lambda_{x}\right)\nu+2\Lambda_{y}^{\dagger}-H_{y}^{\dagger}\right)\cos\beta+\,\left(\left(H_{y}-2\Lambda_{y}\right)\nu-2\Lambda_{x}^{\dagger}+H_{x}^{\dagger}\right)\sin\beta-2\nu\Lambda_{w}\right)\,$, (70a)
$N_{\theta}$ $=4\left(\left(\left(H_{x}^{\dagger}-2\Lambda_{x}^{\dagger}\right)\nu+2\Lambda_{y}-H_{y}\right)\cos\beta+\left(\left(H_{y}^{\dagger}-2\Lambda_{y}^{\dagger}\right)\nu-2\Lambda_{x}+H_{x}\right)\sin\beta-2\nu\Lambda_{w}^{\dagger}\right)\,.$ (70b)
@@
$a_{zj}=\frac{3}{64}\,\sin^{2}(\alpha-\beta)\int_{0}^{\infty}\left(P_{z}|_{\delta=\delta_{j}}\operatorname{ch}(\beta\nu)+Q_{z}|_{\delta=\delta_{j}}\operatorname{sh}(\beta\nu)\right)\operatorname{sch}(\pi\nu)\,\mathrm{d}\nu\,$ (71)
for $j\in\{r,\theta\}$, where $\delta_{r}=0$ and $\delta_{\theta}=\pi/2$. Here,
$P_{z}$ $=4\left(2\Lambda_{z}^{\dagger}-H_{z}^{\dagger}\right)-\left(4\nu^{2}+1\right)\left(\Delta_{x}^{\dagger}\cos\beta+\Delta_{y}^{\dagger}\sin\beta+\Delta_{w}^{\dagger}\right),$ (72a)
$Q_{z}$ $=4\left(2\Lambda_{z}-H_{z}\right)\,-\left(4\nu^{2}+1\right)\left(\Delta_{x}\cos\beta+\Delta_{y}\sin\beta+\Delta_{w}\right).$ (72b)
@@
Conversely, the components $a_{\theta z}$ and $a_{z\theta}$ [panels (b) and (d) in Fig. 5] decrease monotonically with increasing $\beta/\alpha$, reaching their maximum values at $\beta=0$ and vanishing at $\beta=\alpha$. To leading order, no coupling occurs for $\alpha=0$ or $\alpha=\pi/2$.
@@ V.2 Hydrodynamic rotational mobilities
$M_{z}$ $=\left(4\nu\left(H_{x}-2\Lambda_{x}\right)-4\Lambda_{y}^{\dagger}+\left(4\nu^{2}+3\right)H_{y}^{\dagger}\right)\cos\beta+\left(4\nu\left(H_{y}-2\Lambda_{y}\right)+4\Lambda_{x}^{\dagger}-\left(4\nu^{2}+3\right)H_{x}^{\dagger}\right)\sin\beta\,$, (74a)
$N_{z}$ $=\left(4\nu\left(H_{x}^{\dagger}-2\Lambda_{x}^{\dagger}\right)-4\Lambda_{y}+\left(4\nu^{2}+3\right)H_{y}\right)\cos\beta+\left(4\nu\left(H_{y}^{\dagger}-2\Lambda_{y}^{\dagger}\right)+4\Lambda_{x}-\left(4\nu^{2}+3\right)H_{x}\right)\sin\beta\,$, (74b)
@@
$b_{ij}=\frac{1}{16}\,\sin^{3}(\alpha-\beta)\int_{0}^{\infty}\left(P_{i}|_{\delta=\delta_{j}}\operatorname{ch}(\beta\nu)+Q_{i}|_{\delta=\delta_{j}}\operatorname{sh}(\beta\nu)\right)\operatorname{sch}(\pi\nu)\,\mathrm{d}\nu\,$, (75)
@@
$P_{r}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}^{\dagger}\sin\beta-\Delta_{y}^{\dagger}\cos\beta\right)+4\nu\left(2\Lambda_{z}-H_{z}\right)\,$, (76a)
$Q_{r}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}\sin\beta-\Delta_{y}\cos\beta\right)+4\nu\left(2\Lambda_{z}^{\dagger}-H_{z}^{\dagger}\right)\,$, (76b)
@@
$P_{\theta}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}^{\dagger}\cos\beta+\Delta_{y}^{\dagger}\sin\beta\right)+4\Lambda_{z}^{\dagger}-\left(4\nu^{2}+3\right)H_{z}^{\dagger}\,$, (77a)
$Q_{\theta}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}\cos\beta+\Delta_{y}\sin\beta\right)\,+4\Lambda_{z}-\left(4\nu^{2}+3\right)H_{z}\,.$ (77b)

### v2

@@ IV.2 Transverse torque
These results yield expressions for the velocity field that are in full agreement with those reported by Blake and Chwang [61] after conversion from cylindrical to Cartesian coordinates.
@@ V Hydrodynamic mobilities
The solution for the flow field derived above can be used to evaluate how the confining wedge influences the motion of a small spherical colloidal particle of radius $a$ subjected to an external torque. Due to hydrodynamic interactions with the no-slip boundaries, the particle will in general be displaced, and its rotational velocity will be affected. The leading-order term in the hydrodynamic coupling mobility is obtained by evaluating the image solution for the induced flow field at the position of the particle. This procedure yields an expression for the mobilities in terms of the small parameter $\epsilon=a/d$, where $d=\rho\sin(\alpha-\beta)$ and $a$ is the particle radius. Changes in rotational mobility when compared to the situation in an unbounded fluid are obtained by evaluating half the vorticity of the flow resulting from the image solution at the position of the rotating particle.
@@
The translational and rotational velocities resulting from the hydrodynamic interactions with the boundaries are related to the hydrodynamic torque $\bm{L}$ acting on the particle through
@@
$\bm{V}=\left.\bm{v}(\bm{r})-\bm{v}^{\infty}(\bm{r})\right|_{\bm{r}=\bm{r}_{0}}\,,\qquad\bm{\Omega}=\left.\mu_{0}^{r}\bm{L}+\tfrac{1}{2}\,\bm{\nabla}\times\left(\bm{v}(\bm{r})-\bm{v}^{\infty}(\bm{r})\right)\right|_{\bm{r}=\bm{r}_{0}}\,.$ (65)
Here, $\mu_{0}^{r}=1/(8\pi\eta a^{3})$ denotes the bulk hydrodynamic rotational mobility in an unbounded fluid.
Written in a more general form, the grand mobility tensor relates the translational and angular velocities to the forces and torques acting on the particle via
$\begin{pmatrix}\bm{V}\\ \bm{\Omega}\end{pmatrix}=\begin{pmatrix}\bm{M}^{tt}&\bm{M}^{tr}\\ \bm{M}^{rt}&\bm{M}^{rr}\end{pmatrix}\cdot\begin{pmatrix}\bm{F}\\ \bm{L}\end{pmatrix}$. (66)
Vice versa, the grand resistance tensor is defined via
$\begin{pmatrix}\bm{F}\\ \bm{L}\end{pmatrix}=-\begin{pmatrix}\bm{R}^{tt}&\bm{R}^{tr}\\ \bm{R}^{rt}&\bm{R}^{rr}\end{pmatrix}\cdot\begin{pmatrix}\bm{V}\\ \bm{\Omega}\end{pmatrix}$. (67)
We express the elements of the mobility tensor in the point-particle approximation as
$\bm{M}^{tt}=\mu_{0}^{t}\left(\bm{I}-\epsilon\bm{f}^{tt}\right),\quad\bm{M}^{tr}=-\mu_{0}^{c}\,\epsilon^{2}\bm{f}^{tr}\,,\quad\bm{M}^{rr}=\mu_{0}^{r}\left(\bm{I}-\epsilon^{3}\bm{f}^{rr}\right),$ (68)
where $\bm{I}$ denotes the identity matrix and $\bm{M}^{rt}$ is the transpose of $\bm{M}^{tr}$, such that $M_{ij}^{rt}=M_{ji}^{tr}$. Moreover, $\mu_{0}^{t}=1/(6\pi\eta a)$ denotes the hydrodynamic translational mobility in a bulk fluid. $\mu_{0}^{c}=1/(6\pi\eta a^{2})$ is used to scale the mobility coupling coefficients. It follows that, to leading order, the translational mobilities scale as $\epsilon$, the coupling mobilities as $\epsilon^{2}$, while the corrections to the rotational mobilities scale as $\epsilon^{3}$.
Since $\bm{M}$ and $\bm{R}$ are inverses of each other, the elements of $\bm{R}$ can be expressed, to leading order in $\epsilon$, as
$\bm{R}^{tt}=\gamma_{0}^{t}\left(\bm{I}+\epsilon\bm{f}^{tt}\right),\quad\bm{R}^{tr}=\gamma_{0}^{c}\,\epsilon^{2}\bm{f}^{tr}\,,\quad\bm{R}^{rr}=\gamma_{0}^{r}\left(\bm{I}+\epsilon^{3}\bm{f}^{rr}\right).$ (69)
$\bm{R}^{rt}$ is the transpose of $\bm{R}^{tr}$. Here, $\gamma_{0}^{t}=1/\mu_{0}^{t}$, $\gamma_{0}^{r}=1/\mu_{0}^{r}$, and $\gamma_{0}=8\pi\eta a^{2}$ are used to scale the resistance coefficients. For rotational–translational coupling in the hydrodynamic mobilities the scaling is $1/(6\pi\eta a^{2})$, whereas for the corresponding resistance it must be $8\pi\eta a^{2}$ to ensure that the grand mobility and resistance tensors are exact inverses of one another.
Accordingly, when a torque is exerted on the particle, the velocities read
$\bm{V}=-\mu_{0}^{c}\,\epsilon^{2}\bm{f}^{tr}\cdot\bm{L}\,,\qquad\bm{\Omega}=\mu_{0}^{r}\left(\bm{I}-\epsilon^{3}\bm{f}^{rr}\right)\cdot\bm{L}\,$, (70)
where
$\bm{f}^{tr}=\begin{pmatrix}0&0&a_{rz}\\ 0&0&a_{\theta z}\\ a_{zr}&a_{z\theta}&0\end{pmatrix},\qquad\bm{f}^{rr}=\begin{pmatrix}b_{rr}&b_{r\theta}&0\\ b_{\theta r}&b_{\theta\theta}&0\\ 0&0&b_{zz}\end{pmatrix}$ (71)
@@
are dimensionless matrices, corresponding to the rotational–translational coupling mobilities and purely rotational mobilities, respectively. By the symmetry of the mobility tensor, it follows that $b_{r\theta}=b_{\theta r}$. Accordingly, we are left with determining eight components that fully characterize the response to the applied torque. We note that the matrix $\bm{C}$ defined in Eq. (5.2) of Sano and Hasimoto [10] coincides with $\bm{f}^{rt}$, the transpose of $\bm{f}^{tr}$. Accordingly, we have verified that $t_{ij}=a_{ji}$, where $t_{ij}$ denote the elements of $\bm{C}$ as defined in Eq. (5.3) of Sano and Hasimoto [10].
The following results are derived using Eq. (65) together with the solution for the flow field as given by Eq. (52).
@@ V.1 Hydrodynamic coupling mobilities
Figure 5: Variations of the leading-order corrections to the different components of the hydrodynamic coupling mobilities as functions of $\beta/\alpha$ for various values of the semi–opening angle $\alpha$ of the wedge. To leading order, the corrections vanish for $\alpha=\pi/2$ in panels (a) and (c), and for both $\alpha=0$ and $\alpha=\pi/2$ in panels (b) and (d). Some curves are omitted to avoid overcrowding the figure.
@@
$a_{iz}=\frac{3}{64}\,\sin^{2}(\alpha-\beta)\int_{0}^{\infty}\left(M_{i}\operatorname{ch}(\beta\nu)+N_{i}\operatorname{sh}(\beta\nu)\right)\operatorname{sch}(\pi\nu)\,\mathrm{d}\nu\,$, (72)
@@
$M_{r}$ $=\left(12\Lambda_{x}^{\dagger}-\left(4\nu^{2}+7\right)\mathrm{H}_{x}^{\dagger}\right)\cos\beta+\left(12\Lambda_{y}^{\dagger}-\left(4\nu^{2}+7\right)\mathrm{H}_{y}^{\dagger}\right)\sin\beta+4\Lambda_{w}^{\dagger}\,$, (73a)
$N_{r}$ $=\left(12\Lambda_{x}-\left(4\nu^{2}+7\right)\mathrm{H}_{x}\right)\cos\beta+\left(12\Lambda_{y}-\left(4\nu^{2}+7\right)\mathrm{H}_{y}\right)\sin\beta+4\Lambda_{w}\,$, (73b)
@@
$M_{\theta}$ $=4\left(\left(\left(\mathrm{H}_{x}-2\Lambda_{x}\right)\nu+2\Lambda_{y}^{\dagger}-\mathrm{H}_{y}^{\dagger}\right)\cos\beta+\,\left(\left(\mathrm{H}_{y}-2\Lambda_{y}\right)\nu-2\Lambda_{x}^{\dagger}+\mathrm{H}_{x}^{\dagger}\right)\sin\beta-2\nu\Lambda_{w}\right)\,$, (74a)
$N_{\theta}$ $=4\left(\left(\left(\mathrm{H}_{x}^{\dagger}-2\Lambda_{x}^{\dagger}\right)\nu+2\Lambda_{y}-\mathrm{H}_{y}\right)\cos\beta+\left(\left(\mathrm{H}_{y}^{\dagger}-2\Lambda_{y}^{\dagger}\right)\nu-2\Lambda_{x}+\mathrm{H}_{x}\right)\sin\beta-2\nu\Lambda_{w}^{\dagger}\right)\,.$ (74b)
@@
$a_{zj}=\frac{3}{64}\,\sin^{2}(\alpha-\beta)\int_{0}^{\infty}\left(P_{zj}\operatorname{ch}(\beta\nu)+Q_{zj}\operatorname{sh}(\beta\nu)\right)\operatorname{sch}(\pi\nu)\,\mathrm{d}\nu\,$, (75)
for $j\in\{r,\theta\}$, where
$P_{zj}$ $=4\left(2\Lambda_{z}^{\dagger}-\mathrm{H}_{z}^{\dagger}\right)-\left(4\nu^{2}+1\right)\left(\Delta_{x}^{\dagger}\cos\beta+\Delta_{y}^{\dagger}\sin\beta+\Delta_{w}^{\dagger}\right)\Big|_{\delta=0\text{ for }j=r\text{ and }\delta=\frac{\pi}{2}\text{ for }j=\theta\,}$, (76a)
$Q_{zj}$ $=4\left(2\Lambda_{z}-\mathrm{H}_{z}\right)\,-\left(4\nu^{2}+1\right)\left(\Delta_{x}\cos\beta+\Delta_{y}\sin\beta+\Delta_{w}\right)\Big|_{\delta=0\text{ for }j=r\text{ and }\delta=\frac{\pi}{2}\text{ for }j=\theta\,}.$ (76b)
@@
Conversely, the components $a_{\theta z}$ and $a_{z\theta}$ [panels (b) and (d) in Fig. 5] decrease monotonically with increasing $\beta/\alpha$, reaching their maximum values at $\beta=0$ and vanishing at $\beta=\alpha$. It is worth noting that, to leading order, no coupling occurs for $\alpha=0$ or $\alpha=\pi/2$. To quantify the translational motion induced by the rotating particle in these cases, one must include the Laplacian term in Faxén’s law in Eq. (65), which introduces a correction that scales as $\epsilon^{4}$.
To better visualize the meaning of these mobility coefficients, we consider a simple set-up in which a particle is located on the bisector of the wedge and is subject to a torque. This corresponds to setting $\beta=0$, so that the radial and azimuthal directions coincide with the $x$ and $y$ directions, respectively. For a positive axial torque applied along the $z$ direction, the particle undergoes translational motion along the negative $y$ direction, indicated by $a_{rz}=0$ and $a_{\theta z}>0$. The magnitude of the translational velocity depends on the semi–opening angle $\alpha$, and there exists an intermediate angle at which this magnitude is maximal. When the particle is subject to a radial torque, it remains stationary, as expected, which is mathematically corroborated by the fact that $a_{zr}=0$ in this case. For an azimuthal torque applied in the positive direction, the particle undergoes translation along the positive $z$ direction, since $a_{z\theta}<0$. Likewise, there exists an optimal semi–opening angle for which the magnitude of the induced translational velocity is maximal.
It is worth noting that the observed behavior is counterintuitive in that the direction of motion induced by an externally applied torque is opposite to what one might expect. A similar effect has been reported for motion near a fluid–fluid interface [68], where the leading correction, scaling as $\epsilon^{2}$, is directed opposite to intuitive expectations and opposite to that for a sphere near a solid flat wall, for which the leading-order contribution instead scales as $\epsilon^{4}$.
@@ V.2 Hydrodynamic rotational mobilities
$M_{z}$ $=\left(4\nu\left(\mathrm{H}_{x}-2\Lambda_{x}\right)-4\Lambda_{y}^{\dagger}+\left(4\nu^{2}+3\right)\mathrm{H}_{y}^{\dagger}\right)\cos\beta+\left(4\nu\left(\mathrm{H}_{y}-2\Lambda_{y}\right)+4\Lambda_{x}^{\dagger}-\left(4\nu^{2}+3\right)\mathrm{H}_{x}^{\dagger}\right)\sin\beta\,$, (78a)
$N_{z}$ $=\left(4\nu\left(\mathrm{H}_{x}^{\dagger}-2\Lambda_{x}^{\dagger}\right)-4\Lambda_{y}+\left(4\nu^{2}+3\right)\mathrm{H}_{y}\right)\cos\beta+\left(4\nu\left(\mathrm{H}_{y}^{\dagger}-2\Lambda_{y}^{\dagger}\right)+4\Lambda_{x}-\left(4\nu^{2}+3\right)\mathrm{H}_{x}\right)\sin\beta\,$, (78b)
@@
$b_{ij}=\frac{1}{16}\,\sin^{3}(\alpha-\beta)\int_{0}^{\infty}\left(P_{ij}\operatorname{ch}(\beta\nu)+Q_{ij}\operatorname{sh}(\beta\nu)\right)\operatorname{sch}(\pi\nu)\,\mathrm{d}\nu\,$, (79)
@@
$P_{rj}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}^{\dagger}\sin\beta-\Delta_{y}^{\dagger}\cos\beta\right)+4\nu\left(2\Lambda_{z}-\mathrm{H}_{z}\right)\Big|_{\delta=0\text{ for }j=r\text{ and }\delta=\frac{\pi}{2}\text{ for }j=\theta\,}\,$, (80a)
$Q_{rj}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}\sin\beta-\Delta_{y}\cos\beta\right)+4\nu\left(2\Lambda_{z}^{\dagger}-\mathrm{H}_{z}^{\dagger}\right)\Big|_{\delta=0\text{ for }j=r\text{ and }\delta=\frac{\pi}{2}\text{ for }j=\theta\,}\,$, (80b)
@@
$P_{\theta j}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}^{\dagger}\cos\beta+\Delta_{y}^{\dagger}\sin\beta\right)+4\Lambda_{z}^{\dagger}-\left(4\nu^{2}+3\right)\mathrm{H}_{z}^{\dagger}\penalty 10000\ \Big|_{\delta=0\text{ for }j=r\text{ and }\delta=\frac{\pi}{2}\text{ for }j=\theta\,}\,$, (81a)
$Q_{\theta j}$ $=\left(4\nu^{2}+1\right)\left(\Delta_{x}\cos\beta+\Delta_{y}\sin\beta\right)\,+4\Lambda_{z}-\left(4\nu^{2}+3\right)\mathrm{H}_{z}\penalty 10000\ \Big|_{\delta=0\text{ for }j=r\text{ and }\delta=\frac{\pi}{2}\text{ for }j=\theta\,}\,.$ (81b)
@@
To gain further insight into the meaning of these mobility coefficients, we again consider, as a simple example, a particle located on the wedge bisector, so that $\beta=0$. Unlike the coupling mobility, which is absent in a bulk fluid and for which the wedge geometry induces translation through hydrodynamic coupling, the effect of the wedge on rotational motion merely tends to slow down the particle rotation. Moreover, since the correction scales as $\epsilon^{3}$, this effect is even weaker. For a radial torque, the particle undergoes rotation only along the radial direction to this order, because $b_{\theta r}=0$ in this case and $b_{zr}=0$. Smaller opening angles of the wedge lead to larger corrections and therefore to slower rotations; see Fig. 6(a). For an azimuthal torque, we again expect purely azimuthal rotation to this order, because $b_{r\theta}=0$ in this case and $b_{z\theta}=0$. Yet, the dependence on the opening angle of the wedge here is non-monotonic, with a maximal correction at an intermediate angle and a minimum reached in the limit $\alpha=0$; see Fig. 6(b). Finally, for an axial torque, the dependence on the wedge angle is also non-monotonic, exhibiting a minimum in the planar-boundary limit corresponding to $\alpha=\pi/2$; see Fig. 6(c). There is no coupling to the other directional components to the considered order. Overall, the dynamics is strongly dependent on the geometry. As we have indicated, accurate predictions of the mobilities require a systematic quantification.